OpenStudy (anonymous):

find 3 consecutive integers whose sum is 114

hartnn (hartnn):

let the middle number be x so three numbers are x-1,x,x+1 ok? sum = 114 so x-1+x+x+1=114 so 3x=114 so can u find x now? that will be your middle number.

OpenStudy (anonymous):

somebody else told me i could do it this way: 1st number=x, 2nd number=x+2, 3rd number=x+4. So it would be x+x+4+x+2=114, or 3x+6=114. So its 3x=108 and x would be the first number right? Are both ways correct?

hartnn (hartnn):

but they are consecutive , not even (or odd) consecutive means difference = 1 so,x-1,x,x+1 even(or odd) means difference=2 so x,x+2,x+4. ok?

OpenStudy (anonymous):

oh ok. I understand now. Thank you. :)

hartnn (hartnn):

welcome :)